what is the value of x
-3 1/2 +2 3/4x = 5 7/8

Found 2 solutions by josgarithmetic, richwmiller:
Answer by josgarithmetic(39799)

Change all mixed numbers into improper fractions and multiply left and right sides of the equation by 8, and solve....
How you mean the x, is unclear. As written, cannot tell if it is numerator or denominator. Adjust LCD however fits.

Answer by richwmiller(17219)

-3 1/2 +2 3/4x = 5 7/8
common denominator
-3 4/8 +2 6/8x = 5 7/8
2 6/8x = 5 7/8+3 4/8
2 6/8x = 8 7/8+ 4/8
2 6/8x = 8 11/8
now make improper fractions
22x/8=75/8
22x=75
x=75/22